OpenStudy (anonymous):

Can someone please explain in the simplest way how to solve this? Michelle rode her bicycle from her house to school at an average speed of 8 miles per hour. Later that daym she rode school back home along the same route at an average speed of 12 miles per hour. If the round trip took her 1 hour, how many miles is the round trip?

OpenStudy (misty1212):

Hi!!

OpenStudy (misty1212):

don't you hate word problems?

OpenStudy (misty1212):

we can do it though for sure

OpenStudy (anonymous):

ok, thanks!

OpenStudy (misty1212):

we are gong to use distance equals rate time time time, or time is distance divide by rate

OpenStudy (misty1212):

we don't know what the distance is, so lets call it D

OpenStudy (misty1212):

time going it \[\frac{D}{8}\]right?

OpenStudy (anonymous):

yes..

OpenStudy (misty1212):

and time returning is?

OpenStudy (anonymous):

D/12

OpenStudy (misty1212):

right so total time is \[\frac{D}{8}+\frac{D}{12}\] which your are told is 1

OpenStudy (misty1212):

solve \[\frac{D}{8}+\frac{D}{12}=1\] for \(D\) and you are done

OpenStudy (misty1212):

do you know how to do that?

OpenStudy (misty1212):

oh btw they ask for the round trip, so once you find D the answer is actually 2D

OpenStudy (anonymous):

Do we need to find a common denominator first?

OpenStudy (misty1212):

you can to that and then add probably be easier to multiply both sides by the least common multiple of 8 and 12 which is 24 that will clear the annoying fractions

OpenStudy (anonymous):

I did it the first way because it was easier for me, I got D = 4.8

OpenStudy (misty1212):

i didn't do it, let me check

OpenStudy (anonymous):

And then 2*4.8 = 9.6. Thank you so much for helping!

OpenStudy (misty1212):

if you multiply by 24 you get \[3 d+2 d=24\\ 5D=24\\ D=4.8\] yeah you are righ t

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

that was extremely helpful. thank you again. :)

OpenStudy (misty1212):

you are quite welcome